# Application of perturbation theory to finding vibrational frequencies of a spheroid

### v1

@@ I Introduction
In case of elastically isotropic nanocrystals of spherical shape, the frequencies of free vibrations can be found analytically. The solution, involving only trigonometric functions, was obtained in the last quarter of XIX century [11, 12, 13].
For more complex nanocrystal shapes, there have been developed numerical procedures of finding the frequencies of free-standing vibrations [14, 15, 16]. However, it is quite desirable to extend the analytical treatment beyond the case of a sphere. In particular, this would provide a natural way to check the numerical results.
A slightly imperfect sphere can often be described as an ellipsoid. A uniform uniaxial deformation of a sphere yields a spheroid. When this deformation is small, a parameter of non-sphericity can be introduced, and corrections to the vibrational frequencies, linear in this parameter, can be sought. Then a more complex shape of a three-axial ellipsoid can be treated as a superposition of uniaxial deformations along perpendicular axes.
@@
The method seems to be bulletproof in case of the rigid boundary conditions, when the vector field of the displacement is set to be zero at the boundary of the spheroid. However, in case of the free boundary conditions, when traction is set to be zero at the boundary, it is not evident that the stress-free sphere in the transformed coordinates would adequately represent a stress-free spheroid in the initial coordinate system. As the formulation of the method is essentially independent of the particular form of boundary conditions, here we will apply it to both cases and compare the outcomes to the results of numerical solutions.
@@ II Spheroidal vibrations of a spherical nanoparticle
for the rigid boundary conditions, $j_{l}(x)$ are the spherical Bessel functions, ${\bf Y}^{l}_{jm}({\bf r}/r)$ are the vector spherical harmonics [33], the wave numbers $q_{n}$ and $Q_{n}$ are related to the five-fold degenerate mode frequency $\omega_{n}$ through $q_{n}=\omega_{n}/c_{l}$, $Q_{n}=\omega_{n}/c_{t}$, and the frequency is determined as the $(n+1)$-th root of the transcendental equation [31]
@@ III Shape anisotropy
Figure 1: (a): Dimensionless vibration frequencies ($q_{0}R$) of the fundamental spheroidal mode with $j=2$ and $|m|=0,1,2$ as functions of the spheroid aspect ratio, $c/b$, for a restrained CdSe nanocrystal of mean radius $R=c^{1/3}\,b^{2/3}$. The results of the perturbation theory (solid lines) are compared to numerical calculations (lines with symbols). The rigid boundary conditions are used in both cases. (b): Same as (a) but for a silver nanocrystal.
@@ IV Numerical calculations
To test the validity of the perturbation theory approximation, we calculated the vibrations with the free boundary condition using the Rayleigh-Ritz approach as described in Ref. 14. It relies on the choice of products of powers of the Cartesian coordinates ($x^{l}y^{m}z^{n}$) as a basis for expansion of the displacement. This basis was modified for the rigid boundary condition to satisfy the vanishing displacement at the surface of the spheroid by using $x^{l}y^{m}z^{n}\times(1-\frac{x^{2}}{b^{2}}-\frac{y^{2}}{b^{2}}-\frac{z^{2}}{c^{2}})$ functions. These are linear combinations of $x^{l^{\prime}}y^{m^{\prime}}z^{n^{\prime}}$ functions. The integrals of products of the new basis functions over the volume of the spheroid can still be expressed analytically. These are the main ingredients of the numerical calculations. Therefore, the method is rather straightforward to implement and the benefits of the $x^{l}y^{m}z^{n}$ basis functions are maintained.
@@
The validity of the calculations was checked by comparing the obtained frequencies with those of the sphere as determined from Eq. (8) or Eq. (9) but also with frequencies obtained using the finite element method for a few selected spheroids.
@@ V Results and discussion
In Fig. 1 we compare the results of the perturbation theory for CdSe and silver spheroids with mean radii $R=c^{1/3}\,b^{2/3}$ and the rigid boundary conditions with the results of the numerical calculations of vibrational eigenfrequencies. We use the following values of sound velocities: $c_{l}(\mbox{CdSe})=3.7\cdot 10^{5}$ cm/s, $c_{t}(\mbox{CdSe})=1.54\cdot 10^{5}$ cm/s, $c_{l}(\mbox{Ag})=3.747\cdot 10^{5}$ cm/s, $c_{t}(\mbox{Ag})=1.74\cdot 10^{5}$ cm/s and restrict our consideration by the fundamental spheroidal mode with $j=2$ ($_{0}S_{2}^{m}$).
One can see that in both cases of CdSe and Ag nanocrystals the two solutions match in the vicinity of the unit aspect ratio ($c/b=1$). Interestingly, this vicinity is substantially wider in case of silver nanocrystals as compared to CdSe nanocrystals.
@@ VI Conclusions
In this work we have applied the perturbation theory of the boundary conditions in order to find vibrational frequencies of a spheroid from these of a sphere. The perturbation procedure is based on a trick whereby the coordinate change transforms a spheroidal surface to a spherical one of the same enclosed volume. Intuitively, we expect that the boundary condition at the surface, whether it corresponds to a restrained or a stress-free one, formulated for the transformed (spherical) surface, would adequately represent the physical boundary conditions at the spheroidal one as well. However, mathematical requirements imposed on this transformation should be different in the two cases, and, for the later one, the preservation of the angles should be required along with the fact that the Jacobian of the transformation is equal to one (which is equivalent to the preservation of the volume). Since such a transformation apparently does not exist, Migdal’s method fails in case of the stress-free spheroid. However, for the rigid boundary conditions, the perturbation theory works well and provides analytical expression for the splitting of the spheroidal vibrational mode, induced by the symmetry reduction.

### v2

@@ I Introduction
Comparison with atomistic calculations has revealed that the frequencies of fundamental Raman-active vibrational modes can be found within the approximation of elastic continuum for nanocrystals containing as few as 20 atoms [11]. In case of elastically isotropic nanocrystals of spherical shape, the frequencies of free vibrations can be found analytically. The solution, involving only trigonometric functions, was obtained in the last quarter of XIX century [12, 13, 14].
For more complex nanocrystal shapes, numerical procedures of finding the frequencies of free-standing vibrations have been developed [15, 16, 17]. However, it is quite desirable to extend the analytical treatment beyond the case of a sphere. In particular, this would provide a natural way to check the numerical results.
A slightly imperfect sphere can often be described as an ellipsoid. A uniform uniaxial deformation of a sphere yields a spheroid. When this deformation is small, a parameter of non-sphericity can be introduced, and corrections to the vibrational frequencies, linear in this parameter, can be sought. Then a more complex shape of a three-axial ellipsoid can be treated as a superposition of uniaxial deformations along perpendicular axes. Vibrational modes of spheroidal (or almost spheroidal) nanoparticles are also of interest in their own right, as they have been probed in experiments on low-frequency Raman scattering [18, 19].
@@
The method seems to be bulletproof in case of the rigid boundary conditions, when the vector field of the displacement is set to be zero at the boundary of the spheroid. However, in case of the free boundary conditions, when traction is set to be zero at the boundary, it is not evident that the stress-free sphere in the transformed coordinates would adequately represent a stress-free spheroid in the initial coordinate system. As the formulation of the method is essentially independent of the particular form of boundary conditions, here we will apply it to both cases and compare the outcomes to the results of numerical solutions. We note that the rigid boundary conditions are often used when discussing the influence of a matrix surrounding nanocrystals, e.g. in the case of nanocrystals embedded in glass [23]. In practice, the actual frequencies for a nanoparticle embedded in a solid matrix are often intermediate between these yielded by the free and rigid boundary conditions.
@@ II Spheroidal vibrations of a spherical nanoparticle
for the rigid boundary conditions, $j_{l}(x)$ are the spherical Bessel functions, ${\bf Y}^{l}_{jm}({\bf r}/r)$ are the vector spherical harmonics [37], the wave numbers $q_{n}$ and $Q_{n}$ are related to the five-fold degenerate mode frequency $\omega_{n}$ through $q_{n}=\omega_{n}/c_{l}$, $Q_{n}=\omega_{n}/c_{t}$, and the frequency is determined as the $(n+1)$-th root of the transcendental equation [23, 35]
@@ III Shape anisotropy
Figure 1: (a): Dimensionless vibration frequencies ($q_{0}R$) of the fundamental spheroidal mode with $j=2$ and $|m|=0,1,2$ as functions of the spheroid aspect ratio, $c/b$, for a restrained CdSe nanocrystal of mean radius $R=c^{1/3}\,b^{2/3}$. The results of the perturbation theory (solid lines) are compared to numerical calculations (dashed lines). The rigid boundary conditions are used in both cases. Numerical calculations were performed for $R=3$ nm. (b): Same as (a) but for a silver nanocrystal.
@@ IV Numerical calculations
To test the validity of the perturbation theory approximation, we calculated the vibrations with the free boundary condition using the Rayleigh-Ritz approach as described in Ref. 15. It relies on the choice of products of powers of the Cartesian coordinates ($x^{l}y^{m}z^{n}$) as a basis for expansion of the displacement. We used $l+m+n\leq 20$ in this work. No additional restrictions on the basis functions are needed in the case of free boundary conditions, because they emerge as natural boundary conditions from the variational formulation [15, 38, 39]. This basis was modified for the rigid boundary conditions to satisfy the vanishing displacement at the surface of the spheroid by using $x^{l}y^{m}z^{n}\times(1-\frac{x^{2}}{b^{2}}-\frac{y^{2}}{b^{2}}-\frac{z^{2}}{c^{2}})$ functions. These are linear combinations of $x^{l^{\prime}}y^{m^{\prime}}z^{n^{\prime}}$ functions. The integrals of products of the new basis functions over the volume of the spheroid can still be expressed analytically. These are the main ingredients of the numerical calculations. Therefore, the method is rather straightforward to implement and the benefits of the $x^{l}y^{m}z^{n}$ basis functions are maintained.
@@
The validity of the calculations was checked by comparing the obtained frequencies with those of the sphere as determined from Eq. (8) or Eq. (9). The frequency for the sphere is reproduced with a precision of $10^{-8}$ with the Rayleigh-Ritz approach. We also checked the calculations for spheroids with the finite element method for silver spheroids with $b/c=0.6$ and $b/c=1.5$ with the FreeFem++ software [40] using a mesh characteristic length 20 times smaller than the radius of the sphere. The obtained frequencies for the modes of interest matched the Rayleigh-Ritz ones with a precision of $4\times 10^{-4}$.
@@ V Results and discussion
In Fig. 1 we compare the results of the perturbation theory for CdSe and silver spheroids with mean radii $R=c^{1/3}\,b^{2/3}$ and the rigid boundary conditions with the results of the numerical calculations of vibrational eigenfrequencies. We use the following values of sound velocities: $c_{l}(\mbox{CdSe})=3.7\cdot 10^{5}$ cm/s, $c_{t}(\mbox{CdSe})=1.54\cdot 10^{5}$ cm/s, $c_{l}(\mbox{Ag})=3.747\cdot 10^{5}$ cm/s, $c_{t}(\mbox{Ag})=1.74\cdot 10^{5}$ cm/s and restrict our consideration by the fundamental spheroidal mode with $j=2$ ($_{0}S_{2}^{m}$). The numerical calculations were performed for $R=3$ nm.
One can see that in both cases of CdSe and Ag nanocrystals the two solutions match in the vicinity of the unit aspect ratio ($c/b=1$). Interestingly, this vicinity is substantially wider in case of silver nanocrystals as compared to CdSe nanocrystals. This is likely due to a closer proximity of the torsional mode with the total angular momentum $j=3$ ($_{0}T_{3}^{m}$) in the latter case.
@@ VI Conclusions
In this work we have applied the perturbation theory of the boundary conditions in order to find vibrational frequencies of a spheroid from these of a sphere. The perturbation procedure is based on a trick whereby the coordinate change transforms a spheroidal surface to a spherical one of the same enclosed volume. Intuitively, we expect that the boundary condition at the surface, whether it corresponds to a restrained or a stress-free one, formulated for the transformed (spherical) surface, would adequately represent the physical boundary conditions at the spheroidal one as well. However, mathematical requirements imposed on this transformation should be different in the two cases, and, for the latter one, the preservation of the angles should be required along with the fact that the Jacobian of the transformation is equal to one (which is equivalent to the preservation of the volume). Since such a transformation apparently does not exist, Migdal’s method fails in case of the stress-free spheroid. However, for the rigid boundary conditions, the perturbation theory works well and provides analytical expression for the splitting of the spheroidal vibrational mode, induced by the symmetry reduction.
@@ Acknowledgements.
Data availability
The data that support the findings of this article are not publicly available. The data are available from the authors upon reasonable request.